# The Indian Pulsar Timing Array Data Release 2: II. Customised Single-Pulsar Noise Analysis and Noise Budget

### v1

@@ Acknowledgements
ARN DMN FCN SWdet
Pulsars $T_{\text{span}}$ (yr) $\log_{10}A$ $\gamma$ $N_{\text{harm}}$ $\log_{10}A$ $\gamma$ $N_{\text{harm}}$ $\log_{10}A$ $\gamma$ $\chi$ $N_{\text{harm}}$ $n_{\text{earth}}$ $\dot{n}_{\text{earth}}$
J0030++0451 1.76 - - - - - - - - - - $9.62^{+1.26}_{-1.26}$ -
J0034-0534 1.34 - - - - - - - - - - $11.01^{+0.90}_{-0.91}$ -
J0437-4715 6.82 $12.36^{+0.20}_{-0.14}$ $0.37^{+0.48}_{-0.26}$ 81 $3.10^{+0.10}_{-0.11} 1.1$$17^{+0.26}_{-0.28}$ 81 $11.91^{+0.11}_{-0.12}$ $5.03^{+0.71}_{-0.64}$ $0.60^{+0.06}_{-0.06}$ 81 $11.35^{+5.58}_{-7.61}$ -
J0613-0200 7.19 $12.11^{+0.09}_{-0.08}$ $1.58^{+0.25}_{-0.25}$ 86 $13.43^{+0.09}_{-0.08}$ $1.47^{+0.30}_{-0.29}$ 86 - - - - $5.64^{+2.36}_{-2.31}$ -
J0645++5158 2.15 - - - - - - - - - - $7.76^{+6.83}_{-5.40}$ -
J0740++6620 1.93 - - - - - - - - - - $9.76^{+3.77}_{-3.75}$ -
J0751++1807 7.19 - - - - - - - - - - $6.71^{+0.83}_{-0.78}$ -
J0900-3144 1.93 - - - - - - - - - - $9.88^{+6.88}_{-6.72}$ -
J1012++5307 7.10 $12.38^{+0.14}_{-0.19}$ $2.70^{+0.77}_{-0.53}$ 85 - - - - - - - $4.37^{+3.52}_{-2.81}$ -
J1022++1001 7.19 - - - $13.12^{+0.10}_{-0.09}$ $0.40^{+0.28}_{-0.24}$ 86 $16.18^{+0.33}_{-0.19}$ $0.49^{+0.36}_{-0.30}$ $6.68^{+0.24}_{-0.50}$ 60 $11.21^{+1.40}_{-1.40}$ -
J1024-0719 1.08 - - - - - - - - - - $5.95_{-4.30}^{+7.04}$ -
J1125++7819 1.36 - - - - - - - - - - - -
J1455-3330 2.19 - - - - - - - - - - $7.76_{-4.93}^{+5.98}$ -
J1600-3053 5.85 - - - $13.01^{+0.08}_{-0.08}$ $1.45^{+0.25}_{-0.28}$ 70 - - - - $8.67^{+1.92}_{-1.90}$ -
J1614-2230 2.19 - - - - - - - - - - $11.40^{+2.66}_{-2.65}$ -
J1640++2224 2.15 - - - - - - - - - - $11.37^{+6.01}_{-7.19}$ -
J1643-1224 5.90 $11.09^{+0.07}_{-0.11}$ $3.27^{+0.67}_{-0.62}$ 7 $12.07^{+0.05}_{-0.05}$ $3.22^{+0.15}_{-0.14}$ 70 $10.93^{+0.16}_{-0.14}$ $2.75^{+0.86}_{-0.79}$ $0.66^{+0.10}_{-0.09}$ 7 $12.99^{+2.38}_{-2.26}$ -
J1713++0747 4.22 - - - $13.81^{+0.24}_{-1.09}$ $1.44^{+0.94}_{-0.74}$ 50 $16.23^{+0.80}_{-0.72}$ $4.58^{+1.45}_{-1.46}$ $5.31^{+0.99}_{-0.85}$ 50 $1.56^{+2.06}_{-1.11}$ -
J1730-2304 7.20 - - - $13.35^{+0.10}_{-0.11}$ $1.37^{+0.66}_{-0.38}$ 86 - - - - $9.68^{+0.59}_{-0.53}$ -
J1744-1134 2.87 - - - $13.65^{+0.11}_{-0.11}$ $1.08^{+0.36}_{-0.34}$ 34 - - - - $6.26^{+0.60}_{-0.62}$ -
J1857++0943 7.20 - - - $13.47^{+0.10}_{-0.13}$ $2.13^{+0.67}_{-0.52}$ 86 - - - - $6.70^{+3.32}_{-3.29}$ -
J1909-3744 7.20 $12.94^{+0.11}_{-0.11}$ $0.52^{+0.31}_{-0.29}$ 86 $13.54^{+0.09}_{-0.13}$ $1.36^{+0.31}_{-0.27}$ 86 $14.85^{+0.56}_{-0.97}$ $3.01^{+0.87}_{-1.11}$ $3.40^{+1.01}_{-0.50}$ 86 $0.71^{+0.78}_{-0.50}$ $0.85^{+0.17}_{-0.19}$
J1939++2134 5.90 $12.04^{+0.10}_{-0.18}$ $2.21^{+0.39}_{-0.48}$ 59 $12.86^{+0.07}_{-0.06}$ $2.40^{+0.20}_{-0.20}$ 70 $15.01^{+0.26}_{-0.21}$ $2.68^{+0.26}_{-0.25}$ $5.22^{+0.27}_{-0.32}$ 70 $15.67^{+3.12}_{-5.72}$ -
J1944++0907 1.92 - - - $12.86^{+0.16}_{-0.14}$ $3.36^{+0.79}_{-0.64}$ 23 - - - - $9.81^{+6.02}_{5.02}$ -
J2124-3358 5.90 - - - $13.60^{+0.10}_{-0.09}$ $0.31^{+0.26}_{-0.20}$ 70 - - - - $4.80^{+1.92}_{-1.91}$ -
J2145-0750 5.90 - - - $13.59^{+0.11}_{-0.12}$ $0.56^{+0.34}_{-0.31}$ 65 $16.44^{+0.13}_{-0.13}$ $4.03^{+0.81}_{-0.65}$ $6.91^{+0.06}_{-0.14}$ 70 $5.36^{+0.43}_{-0.41}$ -
J2302++4442†† 6.67 - - - - - - - - - - $4.99^{+6.58}_{-3.67}$ -
Table 5: The red noise parameter estimates with 16%-84% credible intervals for the 27 InPTA DR2 pulsars. ARN stands for Achromatic Red Noise, DMN stands for Dispersion Measure Noise, FCN stands for Free Chromatic Noise and SWdet stands for the Deterministic Solar-Wind model where $\dot{n}_{\text{earth}}$ is the first term in the time-domain Taylor-series expansion of the Solar-Wind electron density about a fiducial epoch (which is taken as the starting point of the ToAs for each pulsar). J1125++7819 is a unique case for which the pre-noise residuals are Gaussian, therefore we have only used $\texttt{EFAC}=1.0$ without SWdet for this pulsar. ††PSR J2302++4442 has an apparent $T_{\text{span}}$ of $\sim$$6.67\,\text{yr}$, but its effective $T_{\text{span}}$ is $\sim$$2\,\text{yr}$ due to large data gap (see Figure 8 in Rana+2025).

### v2

@@ Acknowledgements
[…]
Pulsars $T_{\text{span}}$ (yr) Band 3 Band 5 $\log_{10}A$ $\gamma$ $N_{\text{harm}}$ $\log_{10}A$ $\gamma$ $N_{\text{harm}}$ $\log_{10}A$ $\gamma$ $\chi$ $N_{\text{harm}}$ $n_{\text{earth}}$ $\dot{n}_{\text{earth}}$
J0030++0451 1.76 3.91 - - - - - - - - - - - $9.62^{+1.26}_{-1.26}$ -
J0034-0534 1.34 2.17 - - - - - - - - - - - $11.01^{+0.90}_{-0.91}$ -
J0437-4715 6.82 1.44 1.25 $12.36^{+0.20}_{-0.14}$ $0.37^{+0.48}_{-0.26}$ 81 $3.10^{+0.10}_{-0.11} 1.1$$17^{+0.26}_{-0.28}$ 81 $11.91^{+0.11}_{-0.12}$ $5.03^{+0.71}_{-0.64}$ $0.60^{+0.06}_{-0.06}$ 81 $11.35^{+5.58}_{-7.61}$ -
J0613-0200 7.19 1.27 6.12 $12.11^{+0.09}_{-0.08}$ $1.58^{+0.25}_{-0.25}$ 86 $13.43^{+0.09}_{-0.08}$ $1.47^{+0.30}_{-0.29}$ 86 - - - - $5.64^{+2.36}_{-2.31}$ -
J0645++5158 2.15 14.82 24.54 - - - - - - - - - - $7.76^{+6.83}_{-5.40}$ -
J0740++6620 1.93 2.50 5.35 - - - - - - - - - - $9.76^{+3.77}_{-3.75}$ -
J0751++1807 7.19 7.79 10.94 - - - - - - - - - - $6.71^{+0.83}_{-0.78}$ -
J0900-3144 1.93 - 13.95 - - - - - - - - - - $9.88^{+6.88}_{-6.72}$ -
J1012++5307 7.10 6.62 6.77 $12.38^{+0.14}_{-0.19}$ $2.70^{+0.77}_{-0.53}$ 85 - - - - - - - $4.37^{+3.52}_{-2.81}$ -
J1022++1001 7.19 7.03 17.02 - - - $13.12^{+0.10}_{-0.09}$ $0.40^{+0.28}_{-0.24}$ 86 $16.18^{+0.33}_{-0.19}$ $0.49^{+0.36}_{-0.30}$ $6.68^{+0.24}_{-0.50}$ 60 $11.21^{+1.40}_{-1.40}$ -
J1024-0719 1.08 27.70 28.55 - - - - - - - - - - $5.95_{-4.30}^{+7.04}$ -
J1125++7819 1.36 1.71 - - - - - - - - - - - - -
J1455-3330 2.19 24.41 21.20 - - - - - - - - - - $7.76_{-4.93}^{+5.98}$ -
J1600-3053 5.85 4.98 4.35 - - - $13.01^{+0.08}_{-0.08}$ $1.45^{+0.25}_{-0.28}$ 70 - - - - $8.67^{+1.92}_{-1.90}$ -
J1614-2230 2.19 25.35 23.65 - - - - - - - - - - $11.40^{+2.66}_{-2.65}$ -
J1640++2224 2.15 12.37 12.73 - - - - - - - - - - $11.37^{+6.01}_{-7.19}$ -
J1643-1224 5.90 5.08 7.03 $11.09^{+0.07}_{-0.11}$ $3.27^{+0.67}_{-0.62}$ 7 $12.07^{+0.05}_{-0.05}$ $3.22^{+0.15}_{-0.14}$ 70 $10.93^{+0.16}_{-0.14}$ $2.75^{+0.86}_{-0.79}$ $0.66^{+0.10}_{-0.09}$ 7 $12.99^{+2.38}_{-2.26}$ -
J1713++0747 4.22 4.13 4.76 - - - $13.81^{+0.24}_{-1.09}$ $1.44^{+0.94}_{-0.74}$ 50 $16.23^{+0.80}_{-0.72}$ $4.58^{+1.45}_{-1.46}$ $5.31^{+0.99}_{-0.85}$ 50 $1.56^{+2.06}_{-1.11}$ -
J1730-2304 7.20 5.39 17.26 - - - $13.35^{+0.10}_{-0.11}$ $1.37^{+0.66}_{-0.38}$ 86 - - - - $9.68^{+0.59}_{-0.53}$ -
J1744-1134 2.87 1.21 2.32 - - - $13.65^{+0.11}_{-0.11}$ $1.08^{+0.36}_{-0.34}$ 34 - - - - $6.26^{+0.60}_{-0.62}$ -
J1857++0943 7.20 3.50 4.31 - - - $13.47^{+0.10}_{-0.13}$ $2.13^{+0.67}_{-0.52}$ 86 - - - - $6.70^{+3.32}_{-3.29}$ -
J1909-3744 7.20 0.72 1.21 $12.94^{+0.11}_{-0.11}$ $0.52^{+0.31}_{-0.29}$ 86 $13.54^{+0.09}_{-0.13}$ $1.36^{+0.31}_{-0.27}$ 86 $14.85^{+0.56}_{-0.97}$ $3.01^{+0.87}_{-1.11}$ $3.40^{+1.01}_{-0.50}$ 86 $0.71^{+0.78}_{-0.50}$ $0.85^{+0.17}_{-0.19}$
J1939++2134 5.90 0.44 3.55 $12.04^{+0.10}_{-0.18}$ $2.21^{+0.39}_{-0.48}$ 59 $12.86^{+0.07}_{-0.06}$ $2.40^{+0.20}_{-0.20}$ 70 $15.01^{+0.26}_{-0.21}$ $2.68^{+0.26}_{-0.25}$ $5.22^{+0.27}_{-0.32}$ 70 $15.67^{+3.12}_{-5.72}$ -
J1944++0907 1.92 7.78 22.59 - - - $12.86^{+0.16}_{-0.14}$ $3.36^{+0.79}_{-0.64}$ 23 - - - - $9.81^{+6.02}_{5.02}$ -
J2124-3358 5.90 7.94 13.50 - - - $13.60^{+0.10}_{-0.09}$ $0.31^{+0.26}_{-0.20}$ 70 - - - - $4.80^{+1.92}_{-1.91}$ -
J2145-0750 5.90 1.98 6.71 - - - $13.59^{+0.11}_{-0.12}$ $0.56^{+0.34}_{-0.31}$ 65 $16.44^{+0.13}_{-0.13}$ $4.03^{+0.81}_{-0.65}$ $6.91^{+0.06}_{-0.14}$ 70 $5.36^{+0.43}_{-0.41}$ -
J2302++4442†† 6.67 6.77 - - - - - - - - - - - $4.99^{+6.58}_{-3.67}$ -
[…]